# Double Robust Variance Estimation

### v1

@@ Abstract
Doubly robust estimators have gained popularity in the field of causal inference due to their ability to provide consistent point estimates when either an outcome or exposure model is correctly specified. However, the influence function based variance estimator frequently used with doubly robust estimators is only consistent when both the outcome and exposure models are correctly specified. Here, use of M-estimation and the empirical sandwich variance estimator for doubly robust point and variance estimation is demonstrated. Simulation studies illustrate the properties of the influence function based and empirical sandwich variance estimators. Estimators are applied to data from the Improving Pregnancy Outcomes with Progesterone (IPOP) trial to estimate the effect of maternal anemia on birth weight among women with HIV. In the example, birth weights if all women had anemia were estimated to be lower than birth weights if no women had anemia, though estimates were imprecise. Variance estimates were more stable under varying model specifications for the empirical sandwich variance estimator than the influence function based variance estimator.
Keywords causal inference; double robustness; empirical sandwich variance; augmented inverse probability weighting; M-estimation.
@@ 1 Introduction
Two common approaches for estimating the causal effect of a point exposure or treatment with observational studies are methods based on inverse probability of treatment weighting (IPTW) and the g-formula. IPTW estimators (Robins,, 1998; Robins et al.,, 2000; Hernán et al.,, 2000) rely on estimation of the probability of exposure given a set of covariates, i.e., the propensity score. Alternatively, the g-formula (Robins,, 1986) is a direct standardization approach where the outcome is modeled as a function of the exposure and covariates. When estimated nonparametrically, IPTW and g-formula estimators are equivalent (Hernán and Robins,, 2020). However, in high dimensional settings (e.g., settings with continuous covariates), analysts typically resort to parametric models for the exposure and outcome, and the IPTW and g-formula estimators are no longer equivalent due to differing parametric constraints. Both approaches provide consistent estimators of the average causal effect (ACE) if the corresponding model is correctly specified. Because the functional forms of these “nuisance” models are typically unknown, correct specification remains an obstacle to valid effect estimation.
@@
Augmented inverse probability weighted (AIPW) estimators and targeted maximum likelihood estimators (TMLE) combine estimates from the exposure and outcome models in such a way that resulting estimators are “doubly robust.” That is, they are consistent when either the exposure or outcome model is correctly specified, but not necessarily both (Robins et al.,, 2007; Bang and Robins,, 2005; Funk et al.,, 2011; Kang and Schafer,, 2007; Hernán and Robins,, 2020; van der Laan and Rubin,, 2006). In other words, doubly robust methods give the analyst two opportunities for correct model specification. Two common AIPW estimators are an estimator that incorporates estimated propensity scores from the exposure model and pseudo-outcomes from the outcome model (i.e., the classic AIPW) (Hernán and Robins,, 2020; Lunceford and Davidian,, 2004) and a g-formula type estimator where the outcome model is weighted by the IPTW (i.e., the weighted regression AIPW) (Kang and Schafer,, 2007; Robins et al.,, 2007; Vansteelandt and Keiding,, 2011; Gabriel et al.,, 2023). TMLE is a variation of the weighted regression AIPW estimator that instead combines IPTW and the outcome model through a “targeting” model fit using the outcome predictions and weights (van der Laan and Rubin,, 2006). Both the weighted regression AIPW estimator and TMLE respect the parameter space of the estimand, while the classic AIPW can produce estimates outside of the range of the parameter space (Gabriel et al.,, 2023; Gruber and Van Der Laan,, 2012).
@@
In addition to the double robustness property, doubly robust estimators also offer precision advantages over IPTW estimators. When both nuisance models are correctly specified, the three doubly robust estimators previously discussed are semiparametric efficient, having the smallest asymptotic variance of all IPTW estimators (Robins et al.,, 1994; van der Laan and Rubin,, 2006; Gabriel et al.,, 2023). This means that even in settings where the analyst is confident in the specification of one of the models (e.g., the exposure mechanism is well known), doubly robust estimators can offer improved efficiency relative to IPTW estimators, provided that the outcome model is not grossly misspecified. Given these potential precision advantages and that the true specification of both models is typically unknown, doubly robust approaches have gained popularity (Sullivan et al.,, 2022; Josey et al.,, 2023; Lee et al.,, 2019; Zhong et al.,, 2018; Gerdts et al.,, 2022; Chua et al.,, 2022; Hoffman et al.,, 2022; Yazdi et al.,, 2021).
@@
Despite adoption of doubly robust estimators, there is a major limitation with the corresponding influence function based variance estimator commonly used in applications and provided by software. Namely, this variance estimator is only consistent if both the propensity and outcome models are correctly specified (Daniel,, 2014). This limitation has been highlighted previously (Funk et al.,, 2011; Daniel,, 2014), yet the influence function based variance estimator is still commonly used. A systematic review published in October 2023 found that 67% of papers published using TMLE did not indicate how variances were estimated. Of the 33% that did report the method, the majority (73%) relied on the influence function based variance estimator (Smith et al.,, 2023). Widespread application of the influence function based variance estimator is aided by available software, as it is used for confidence interval (CI) construction in the R AIPW (Zhong et al.,, 2021) and tmle (Gruber and Van Der Laan,, 2012) packages as well as the zEpid (Zivich et al., 022a, ) Python package.
The influence function based variance estimator’s reliance on correctly specifying both exposure and outcome models is particularly concerning in instances where researchers understand either the exposure or outcome mechanism well, but not both. If the outcome process is well known but the exposure process is not well understood, leading to correct specification of the outcome model but misspecification of the propensity model, the influence function based variance estimator is inconsistent (Laan and Robins,, 2003; Muñoz and Van Der Laan,, 2012). In this setting, consistent point estimators may be paired with CIs that do not achieve the nominal level of coverage. Alternatively, if the exposure process is well understood but the outcome process is not, leading to correct specification of the propensity model but misspecification of the outcome model, the influence function based variance estimator is conservative (Muñoz and Van Der Laan,, 2012). Precision advantages of doubly robust estimators may be lost with application of a conservative variance estimator. For example, we aim to estimate the effect of maternal anemia on birth weight among women with HIV. While most causes of anemia among women with HIV are well understood, doubly robust estimators may be preferred over IPTW estimators due to efficiency advantages. Modeling the relationship between anemia and birth weight among women with HIV is challenging due to complex dependencies between anemia and other drivers of birth weight. Furthermore, low birth weight may result from two separate pathways - preterm birth or fetal growth restrictions - each of which is a complex syndrome with several causes. For these reasons, the outcome model may be susceptible to misspecification, thus application of the influence function based variance estimator may lead to conservative variance estimates in this study.
@@
Alternative, but less commonly used, variance estimation approaches are themselves doubly robust, providing consistent estimates of the variance when only one of the two nuisance models is correctly specified. Two such approaches are the nonparametric bootstrap and the empirical sandwich variance estimator based on M-estimation theory. The nonparametric bootstrap, while doubly robust (Funk et al.,, 2011), is computationally burdensome. For large datasets, this computational burden may preclude its use. M-estimation (Stefanski and Boos,, 2002) has been proposed for use with doubly robust estimators (Lunceford and Davidian,, 2004; Daniel,, 2014), but we are not aware of prior work highlighting the doubly robust property of the empirical sandwich variance estimator. Furthermore, the recent systematic review (Smith et al.,, 2023) did not mention application of the empirical sandwich variance estimator in any of the papers evaluated. This suggests that the doubly robust property of the empirical sandwich variance estimator is not well known.
@@
This paper demonstrates how M-estimation can be applied for AIPW estimators and TMLE to obtain doubly robust point and variance estimators. In Section 2, three common doubly robust estimators are described along with two approaches for inference: the influence function based variance estimator and M-estimation paired with the empirical sandwich variance estimator. Estimators are applied to data from the Improving Pregnancy Outcomes with Progesterone (IPOP) trial to estimate the average causal effect of anemia on birth weight in Section 3. The doubly robust property of the empirical sandwich variance estimator is demonstrated in a simulation study and is contrasted with the performance of the influence function based variance estimator in Section 4. The benefits and limitations of each variance estimation approach are discussed in Section 5. R code for implementing each of the doubly robust estimators is provided on GitHub.
@@ 2.2 Estimators
Three doubly robust approaches for estimating the $ACE$ are considered: the classic AIPW estimator, the weighted regression AIPW estimator, and TMLE. All three methods combine estimates from a model for the probability of exposure conditional on covariates (propensity model) and a model for the outcome conditional on exposure and covariates (outcome model). The approach for modeling the propensity score is common across the three methods, while there are slight differences in fitting the outcome model. For all three approaches, the propensity score for each participant, $e_{i}=\Pr(X_{i}=1\mid Z_{i})$, is estimated. In practice, often the model $\mbox{logit}(e_{i})=G_{i}^{T}\alpha$ is fit, where $G_{i}=g(Z_{i})$ is a vector of predictors for participant $i$ for some user-specified function $g$ of $Z_{i}$ and $\alpha$ is the vector of regression coefficients from the propensity model. The predicted propensity score for each participant is calculated as $\hat{e}_{i}=e(Z_{i},\hat{\alpha})=\mbox{logit}^{-1}(G_{i}^{T}\hat{\alpha})$ where $\hat{\alpha}$ is the maximum likelihood estimate (MLE) of $\alpha$. The IPTW is then estimated for each participant as $\hat{W}_{i}=X_{i}\hat{e}_{i}^{-1}+(1-X_{i})(1-\hat{e}_{i})^{-1}$.
@@ 2.2.1 Classic AIPW estimator
For the classic AIPW estimator, the $Y\mid Z,X$ model is fit and used to predict pseudo-outcomes for each participant under both exposures. That is, $a_{x}(Z_{i},\hat{\gamma})=\hat{E}(Y_{i}\mid Z_{i},X_{i}=x)$ for $x\in\{0,1\}$, where $\hat{\gamma}$ are the MLEs of the parameters from the assumed outcome model. Then, the classic AIPW estimator for the $ACE$ is
@@ 2.2.2 Weighted regression AIPW estimator
An alternative AIPW estimator incorporates the estimated propensity scores when fitting the outcome regression model. The parameters of the $Y\mid Z,X$ model are estimated with IPTW-weighted maximum likelihood estimation. Pseudo-outcomes are obtained for each participant under both exposures, i.e., $b_{x}(Z_{i},\hat{\beta})=\hat{E}(Y_{i}\mid Z_{i},X_{i}=x)$ for $x\in\{0,1\}$, where $\hat{\beta}$ are the MLEs from the weighted outcome model. The weighted regression AIPW estimator is
@@
where $\hat{\mu}_{WR}^{x}=n^{-1}\sum_{i=1}^{n}b_{x}(Z_{i},\hat{\beta})$ for $x\in\{0,1\}$. The weighted regression AIPW estimator (2) has been evaluated previously (Kang and Schafer,, 2007; Robins et al.,, 2007; Vansteelandt and Keiding,, 2011; Gabriel et al.,, 2023) and is expected to be more stable than (1) when IPTWs are extreme due to bounding within the parameter space (Kang and Schafer,, 2007; Robins et al.,, 2007; Vansteelandt et al.,, 2012).
@@ 2.2.3 TMLE
TMLE is an alternative doubly robust estimator that consists of a two-stage process. In the first stage, $Y\mid Z,X$ is modeled as in Section 2.2.1, and $a_{x}(Z_{i},\hat{\gamma})$ are calculated for $x\in\{0,1\}$. In the second stage, referred to as the targeting stage (van der Laan and Rubin,, 2006), corrections to outcome regression predictions from stage one are made that incorporate the estimated propensity scores. Different methods have been proposed for fitting the targeting model; here, the weighted regression approach is considered, as it is thought to handle random positivity violations better (van der Laan et al.,, 2011). Specifically, the models $\mbox{logit}(Y_{i})=\eta_{x}+\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}$ are fit for $x\in\{0,1\}$. Parameters $\eta_{x}$ are estimated using weighted maximum-likelihood, with weights $(1-X)\hat{W}_{i}$ and $X\hat{W}_{i}$ for the $x=0$ and $x=1$ targeting model, respectively. In the targeting models, logit-transformed pseudo-outcomes ($\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}$) for each participant are included as offset terms, and MLEs $\hat{\eta}_{0}$ and $\hat{\eta}_{1}$ can be thought of as corrections to these pseudo-outcomes (Zivich et al., 022b, ). The stage two pseudo-outcomes are computed as $c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}+\hat{\eta}_{x}]$. The TMLE estimator of $ACE$ is
@@
where $\hat{\mu}_{TMLE}^{x}=n^{-1}\sum_{i=1}^{n}c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})$ for $x\in\{0,1\}$. Note that for continuous outcomes, such as birth weight considered in Section 3, the outcome is scaled prior to implementing TMLE, i.e., $Y_{i}^{*}=(Y_{i}-a)/(b-a)$ is defined for each participant, where $(a,b)$ are the bounds of $Y$. Then, $Y_{i}$ is replaced with $Y_{i}^{*}$ in the steps above, and $c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}+\hat{\eta}_{x}](b-a)+a$ re-scales the outcome back to the original bounds of $Y$.
TMLE methods were originally proposed by van der Laan and Rubin, (2006) and have been extended and applied in a number of settings (Stitelman et al.,, 2012; Balzer et al.,, 2023; van der Laan et al.,, 2018; van der Laan and Gruber,, 2012). When parametric models are used in both stages of estimation, (3) is expected to be similar to (2) (Tran et al.,, 2019).

### v2

@@ Abstract
Doubly robust estimators have gained popularity in the field of causal inference due to their ability to provide consistent point estimates when either an outcome or exposure model is correctly specified. However, for nonrandomized exposures the influence function based variance estimator frequently used with doubly robust estimators of the average causal effect is only consistent when both working models (i.e., outcome and exposure models) are correctly specified. Here, the empirical sandwich variance estimator and the nonparametric bootstrap are demonstrated to be doubly robust variance estimators. That is, they are expected to provide valid estimates of the variance leading to nominal confidence interval coverage when only one working model is correctly specified. Simulation studies illustrate the properties of the influence function based, empirical sandwich, and nonparametric bootstrap variance estimators in the setting where parametric working models are assumed. Estimators are applied to data from the Improving Pregnancy Outcomes with Progesterone (IPOP) study to estimate the effect of maternal anemia on birth weight among women with HIV.
Keywords augmented inverse probability weighting; causal inference; double robustness; empirical sandwich variance; M-estimation.
@@ 1 Introduction
Two common approaches for estimating causal effects of point exposures with observational study data are inverse probability of treatment weighting (IPTW) and g-computation. IPTW estimators rely on estimation of the probability of exposure given a set of covariates, i.e., the propensity score. Alternatively, g-computation is a direct standardization approach where the outcome is modeled as a function of the exposure and covariates. Both approaches provide consistent estimators of the average causal effect (ACE) if the corresponding model is correctly specified, but these estimators are generally not equivalent due to differing parametric constraints imposed by modeling. As the functional forms of these working, or nuisance, models are typically unknown, correct specification remains an obstacle to valid effect estimation.
@@
Augmented inverse probability weighted (AIPW) estimators and targeted maximum likelihood estimators (TMLE) combine estimates from the working models in such a way that resulting estimators are “doubly robust.” That is, they are consistent when either the exposure or outcome model is correctly specified, but not necessarily both (Robins et al.,, 2007; Funk et al.,, 2011; Kang and Schafer,, 2007; van der Laan and Rubin,, 2006). Two common variations of AIPW estimators are an estimator that incorporates estimated propensity scores from the exposure model and pseudo-outcomes from the outcome model (i.e., the classic AIPW) and a g-computation type estimator where the outcome model is weighted by the IPTW (i.e., the weighted regression AIPW). TMLE is a variation of the weighted regression AIPW estimator that instead combines IPTW and the outcome model through a “targeting” model fit using the outcome predictions and weights. While primarily developed with machine learning approaches for estimation of nuisance parameters (van der Laan and Rubin,, 2006), TMLE can also be applied with parametric working models.
@@
Doubly robust estimators also offer precision advantages over IPTW estimators. When both working models are correctly specified, the three doubly robust estimators previously discussed are semiparametric efficient, having the smallest asymptotic variance of all IPTW estimators (Robins et al.,, 1994; van der Laan and Rubin,, 2006; Gabriel et al.,, 2023). Given these potential precision advantages and that the true specification of both models is typically unknown, doubly robust approaches have gained popularity.
@@
The influence function based variance estimator commonly used with doubly robust methods has a major limitation. When estimating the average causal effect from observational data, this variance estimator is only consistent if both working models are correctly specified, i.e., this variance estimator is not doubly robust. This limitation has been highlighted previously (Funk et al.,, 2011; Daniel,, 2014), yet the influence function based variance estimator is still commonly used. A systematic review published in October 2023 found that 67% of papers published using TMLE did not indicate how variances were estimated. Of the 33% that did report the method, the majority (72%) relied on the influence function based variance estimator (Smith et al.,, 2023). Widespread application of the influence function based variance estimator is aided by available software, as it is used for confidence interval (CI) construction in the R AIPW and tmle packages and the zEpid Python package.
When estimating the average causal effect, performance of the influence function based variance estimator depends on which working model is misspecified. When the outcome model is correctly specified but the propensity model is misspecified, the influence function based variance estimator is inconsistent (Muñoz and Van Der Laan,, 2012). In this setting, consistent point estimators may be paired with CIs that do not achieve the nominal level of coverage. If the propensity model is correctly specified but the outcome model is misspecified, the influence function based variance estimator is conservative (Muñoz and Van Der Laan,, 2012). Precision advantages of doubly robust estimators may be lost with application of a conservative variance estimator. These properties of the influence function based variance estimator do not necessarily apply for other estimands and in the randomized setting (see Section 5).
@@
Alternative, but less commonly used, variance estimation approaches are doubly robust, as they are expected to provide valid estimates of the variance leading to nominal CI coverage when only one working model is correctly specified. Two such approaches are the nonparametric bootstrap and the empirical sandwich variance estimator based on M-estimation theory. The doubly robust property of the nonparametric bootstrap has been illustrated previously (Funk et al.,, 2011). M-estimation has been proposed for use with doubly robust estimators (Lunceford and Davidian,, 2004; Daniel,, 2014), but we are not aware of prior work highlighting the doubly robust property of the empirical sandwich variance estimator. The M-estimation approach is equivalent to using the correct influence function with correction terms for estimation of nuisance parameters (Stefanski and Boos,, 2002).
The recent systematic review (Smith et al.,, 2023) did not mention application of the empirical sandwich variance estimator in any reviewed papers. M-estimation is appropriate with finite-dimension parametric working models and is generally not compatible with many machine learning approaches. SuperLearner was used in the majority of studies considered in the systematic review (62%), and the empirical sandwich variance estimator would be inappropriate for these studies. However, logistic regression was used in 20% of studies considered, and an additional 18% of studies did not indicate what methods were used for nuisance parameter estimation. The lack of use of the empirical sandwich variance estimator in these settings suggests that the doubly robust property of the empirical sandwich variance estimator is not well known. Among published studies that reported the variance estimation approach, 21% applied the bootstrap. However, knowledge of its doubly robust property may or may not be widespread.
We aim to estimate the effect of maternal anemia on birth weight among women with HIV. Prior research has shown links between maternal anemia and low birth weight (Azizah et al.,, 2022), and women with HIV experience anemia more frequently than women without HIV (Levine et al.,, 2001). This effect is estimated with data from approximately 800 participants from the Improving Pregnancy Outcomes with Progesterone (IPOP) study. IPOP was a randomized controlled trial conducted in Lusaka, Zambia to evaluate the effectiveness of a therapeutic for prevention of preterm birth and stillbirth among women with HIV. Additional details about the IPOP trial have been published previously (Price et al.,, 2021). Covariates collected during the trial allow for examination of the relationship between anemia and birth weight. Doubly robust estimators may be preferred over IPTW or g-computation approaches due to uncertainty in working model specifications and potential efficiency advantages.
@@
This paper demonstrates how M-estimation and the nonparametric bootstrap can be applied with AIPW estimators and TMLE to obtain doubly robust point and variance estimators with parametric working models. In Section 2, three common doubly robust estimators are described along with three approaches for inference: the influence function based variance estimator, the empirical sandwich variance estimator, and the nonparametric bootstrap. Estimators are applied to IPOP data in Section 3. In Section 4, the doubly robust properties of the empirical sandwich variance estimator and the nonparametric bootstrap are demonstrated in a simulation study and are contrasted with the performance of the influence function based variance estimator. The benefits and limitations of each variance estimation approach are discussed in Section 5.
@@ 2.2 Estimators
Three doubly robust estimators of $ACE$ are considered, each of which combines estimates from the propensity and outcome models. The approach for modeling the propensity score is common across the methods, while there are slight differences in fitting the outcome model. In practice, the propensity score for each participant, $e_{i}=\Pr(X_{i}=1\mid Z_{i})$, is often estimated from the fitted model $\mbox{logit}(e_{i})=G_{i}^{T}\alpha$, where $G_{i}=g(Z_{i})$ is a vector of predictors for participant $i$ for some user-specified function $g$ of $Z_{i}$ and $\alpha$ is the vector of regression coefficients from the propensity model. The predicted propensity score for each participant is calculated as $\widehat{e}_{i}=e(Z_{i},\widehat{\alpha})=\mbox{logit}^{-1}(G_{i}^{T}\hat{\alpha})$ where $\hat{\alpha}$ is the maximum likelihood estimate (MLE) of $\alpha$. The IPTW is then estimated for each participant as $\hat{W}_{i}=X_{i}\hat{e}_{i}^{-1}+(1-X_{i})(1-\hat{e}_{i})^{-1}$.
@@ 2.2.1 Classic AIPW estimator
For the classic AIPW estimator, the $Y\mid Z,X$ model is fit and used to predict pseudo-outcomes for each participant under both exposures. That is, $a_{x}(Z_{i},\hat{\gamma})=\hat{E}(Y_{i}\mid Z_{i},X_{i}=x)$ for $x\in\{0,1\}$, where $\hat{\gamma}$ are the MLEs of the parameters from the assumed outcome model. Then, $ACE$ is estimated by
@@ 2.2.2 Weighted regression AIPW estimator
For the weighted regression AIPW estimator, the parameters of the $Y\mid Z,X$ model are estimated with IPTW-weighted maximum likelihood estimation. Pseudo-outcomes are obtained for each participant under both exposures, i.e., $b_{x}(Z_{i},\hat{\beta})=\hat{E}(Y_{i}\mid Z_{i},X_{i}=x)$ for $x\in\{0,1\}$, where $\hat{\beta}$ are the MLEs from the weighted outcome model. The weighted regression AIPW estimator is
@@
where $\hat{\mu}_{WR}^{x}=n^{-1}\sum_{i=1}^{n}b_{x}(Z_{i},\hat{\beta})$ for $x\in\{0,1\}$. The estimator (2) has been evaluated previously and is expected to be more stable than (1) when IPTWs are extreme due to bounding within the parameter space (Kang and Schafer,, 2007; Robins et al.,, 2007; Vansteelandt et al.,, 2012).
@@ 2.2.3 TMLE
TMLE consists of a two-stage process. In the first stage, $Y\mid Z,X$ is modeled as in Section 2.2.1, and $a_{x}(Z_{i},\hat{\gamma})$ are calculated for $x\in\{0,1\}$. In the second stage, referred to as the targeting stage, corrections to outcome regression predictions from stage one are made that incorporate the estimated propensity scores. Different methods have been proposed for fitting the targeting model; here, the weighted regression approach is considered, as it is thought to handle random positivity violations better (van der Laan et al.,, 2011). First, assuming $Y$ is binary, the models $\mbox{logit}\{\Pr(Y_{i}=1\mid Z_{i},X_{i}=x)\}=\eta_{x}+\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}$ are fit for $x\in\{0,1\}$. Note the only parameter in this model is $\eta_{x}$, as $\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}$ is a fixed offset. Parameters $\eta_{x}$ are estimated using weighted maximum likelihood, with weights $(1-X)\hat{W}_{i}$ and $X\hat{W}_{i}$ for the $x=0$ and $x=1$ models, respectively. The stage two pseudo-outcomes are computed as $c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}+\hat{\eta}_{x}]$. Then, the TMLE estimator of $ACE$ is
@@
where $\hat{\mu}_{TMLE}^{x}=n^{-1}\sum_{i=1}^{n}c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})$ for $x\in\{0,1\}$. For continuous outcomes, a logit model is still recommended for the targeting stage to ensure bounding within the parameter space (Gruber and van der Laan,, 2010; Gruber and Van Der Laan,, 2012). Here, the outcome is scaled prior to implementing TMLE, i.e., $Y_{i}^{*}=(Y_{i}-a)/(b-a)$ is defined, where $(a,b)$ are the bounds of $Y$. Then, $Y_{i}$ is replaced with $Y_{i}^{*}$ in the steps above, i.e., $E(Y_{i}^{*}\mid Z_{i},X_{i}=x)$ is modeled with a logit link, and $c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}+\hat{\eta}_{x}](b-a)+a$ re-scales the outcome back to the original bounds of $Y$.
TMLE methods were developed by van der Laan and Rubin, (2006) and are frequently paired with machine learning techniques for nuisance modeling. When parametric models are used in both stages of estimation, (3) is expected to be similar to (2) (Tran et al.,, 2019).